# Piezoelectric actuation for integrated photonics

### v1

@@ 4.1.2 Wavelength Switchable RSOA-Based External Cavity Laser
There are two frequency tuning schemes supported by the Vernier filter. The first one is wavelength switching by tuning the resonance alignment of the two resonators from one pair to another that is one FSR away, as illustrated in Fig. 25(e). This frequency tuning is discrete but can cover a large range (determined by the FSR) by only tuning a small amount of the resonant frequency of one microring (determined by the FSR difference), which is preferable for piezoelectric actuators. In the experiment, a heater is necessary to initially align the two resonators which are generally far off from each other due to the fabrication error. Following that, the resonance of one resonator can be tuned by its piezoelectric actuator relative to the other. The wavelength switching direction is determined by which microring is actuated. Figure 25(f) shows the emission wavelength under two switching states that are separated by an FSR. The switching speed is tested by applying 8 V 100 kHz square wave and a 7 ns rise time is achieved. The second scheme is the continuous frequency tuning by actuating the two microrings concurrently. As illustrated in Fig. 25(g), a frequency excursion of 1.5 GHz is generated under 100 kHz triangular wave with 11 V. To suppress the nonlinearity, two algorithms are combined to correct the input voltage, including the frequency response inversion [257] and interative signal correction. After linearization, the RMS nonlinearity changes from 15 MHz to 3.8 MHz, as shown in Fig. 25(h).
@@ 4.1.3 High Power and Tunable Extended-DBR Laser
Table 4 summarizes the performances of the piezoelectric tuned hybrid lasers and compares them with the state-of-the-art hybrid or heterogeneous integrated lasers with fixed frequency, as well as lasers that are tuned by the heater or electro-optic effect. We can see most of the lasers with sub-kHz linewidth are achieved by self-injection locking a DFB laser or an ECL with E-DBR to an ultra-high Si3N4 resonator. The frequency noise is primarily limited by the thermal-refractive noise (TRN) of the Si3N4 [253]. A recent work further suppresses noise beyond TRN by locking an E-DBR SIL laser to a Fabry-Pérot cavity via the Pound–Drever–Hall technique [251]. While the PDH signal is currently generated by the off-chip electro-optic modulator, the same signal has been generated on-chip by the piezoelectric actuator that is compatible with the Si3N4 PICs [54]. The thermo-optic tuning has achieved the largest tuning range at the expense of high power consumption and slow speed, and the tuning range decreases as the speed increases. On the other hand, electro-optic tuning based on LiNbO3 PICs presents the highest speed up to 600 MHz. However, the laser linewidth is above kHz even with SIL. A recent study reveals that the charge noise of the LiNbO3 ultimately sets the noise floor of the hybrid laser [268]. As for the piezoelectric tuning, lasers with sub-kHz linewidth, over 1 GHz tuning range, and hundreds of kHz modulation speed have been demonstrated. So far, these experiments are implemented by edge coupling the DFB laser or gain chip to the Si3N4 photonic chip. In the furture, the development heterogeneous integration of the laser with piezoelectric actuators and the photonic wirebonding technique will be of great interest for building a more compact and stable chip-scale laser system [246, 250].
@@ 4.2.1 Voltage Initiation and Tuning of Soliton Microcombs
4.2.2 Fast Soliton Actuation and Locking
@@
The above experiment only demonstrates the stabilization of one comb line that is involved in the feedback loop. To show the stability of the entire comb, the out-of-loop comb needs to be characterized. To measure the out-of-loop beat signal and its phase noise, a modified setup is used as shown in Fig. 29(a). The pump laser’s frequency to generate the soliton microcomb is shifted by 77.0 MHz via a fibre-coupled acousto-optic modulator (AOM). The reason to shift the microcomb’s pump frequency is to cancel out the drift and noise caused by the imbalanced paths in delayed self-homodyne measurement, by detecting the beatnote (77.0 MHz shift) between the pump lines of the microcomb and the EO comb. By down-mixing the 77.0 MHz heterodyne beatnote signal using the same microwave source that drives the AOM, the feedback signal is applied to the laser current such that the pump laser’s frequency is stabilized and the noise in the delayed self-homodyne measurement is removed.
@@
In addition to the quasi-DC tuning, the HBAR modes of the AlN actuator provide a novel functionality for error signal generation using the Pound-Drever-Hall (PDH) technique. Figure 31(a) shows the experimental setup to generate PDH error signals using HBAR modes. The measured response of the AlN actuation, up to 400 MHz, is plotted in the linear frequency scale in Fig. 31(b), showing both cases when the laser is on- and off-resonance. Different modulation frequencies corresponding to different HBAR modes are investigated, which are marked with stars in Fig. 31(b). The PDH error signals modulated at these HBAR frequencies are shown in Fig. 31(c), as well as the corresponding microresonator’s resonance. A microwave source providing 8 dBm RF power is used to modulate the Si3N4 microresonator via AlN actuation. The decrease in error signal contrast at higher HBAR frequency is caused by the finite bandwidth of the optical resonance. This PDH signal can be used to stabilize the laser frequency referenced to an on-chip high optical Q Si3N4 microresonator.
4.2.3 Soliton-Based Parallel LiDAR Engine
The AlN actuator can be a key component for a soliton-based, parallel, frequency-modulated continuous-wave (FMCW) LiDAR [38]. This scheme conventionally relies on scanning the pump laser’s frequency over the soliton existence range , i.e. , with being the resonance frequency, and the boundary of soliton existence detuning range. As a result, the pump frequency chirp is transferred to all soliton comb lines, as shown in Fig. 32(a). The diffractive optics subsequently separates the comb lines spatially, which allows high-speed and parallel acquisition of both velocity and position in each pixel. However, the varying soliton detuning has several limitations. First, the soliton detuning variation leads to changes in soliton spectrum bandwidth and power, which limits the number of usable optical channels. Second, the Raman self-frequency shift [277, 278] causes a varied soliton repetition rate. Third, GHz frequency excursion of the pump laser () requires GHz-wide soliton existence range (), which leads to elevated pump power to several Watts.
@@ 4.2.3 Soliton-Based Parallel LiDAR Engine
In conclusion, integrated piezoelectric actuators that are linear, fast, non-absorptive, bidirectional, and consume ultralow electric power endow integrated soliton microcombs with novel capability in power-critical applications, e.g., in space, data centers and portable atomic clocks, in extreme environments such as cryogenic temperatures, or in emerging applications for coherent LiDAR [38], frequency synthesizers [34] and RF photonics [280, 281]. Therefore, this novel capability is important for tuning and stabilization of soliton microcombs, and may allow soliton injection locking for stabilization and microwave synthesis [274, 282]. While polycrystalline AlN is used in current works, the operation voltage can be reduced by more than 2 times using scandium-doped AlN [283]. In the future, the co-integration of CMOS electronic circuitry on a close-by die paves the way for packaged soliton microcombs with rapid electronic actuation [260].
5 Microwave-Frequency Acousto-Optic Modulation
While piezoelectric actuators have found wide applications in tuning of integrated photonic devices, fast actuation at microwave frequencies will excite acoustic waves that enable efficient acousto-optic modulation (AOM). At the acoustic resonance, integrated AOMs have been demonstrated with efficiency comparable to or even higher than state-of-the-art electro-optic modulators, albeit at the expense of narrow bandwidth [15, 14, 12, 284, 285]. Nevertheless, it has been applied not only in classical applications, but also in quantum regime working as quantum transducers between microwave and light [189, 286]. In the following, we will review recent progress on chip-scale AOMs that can be classified into three categories in terms of the acoustic mode being utilized: surface acoustic wave (SAW), bulk acoustic wave (BAW), and optomechanical crystal (OMC).
@@ 5.1 Surface Acoustic Wave Based AOM
The surface acoustic wave excited by interdigital transducer (IDT) has been one of the traditional ways of launching acoustic waves on a chip [287], and has become the key technology across a wide range of research areas, such as microfluidics [288], biosensors [289], radio-frequency filters and delay lines [290, 291], and quantum acoustics [292]. As nano-fabrication advances, the ability to create sub-micron electrode fingers has extended the frequency of the acoustic wave to beyond 10 GHz [293, 133]. By interacting with an optical cavity, the stress and strain field of the acoustic wave can be used to modulate the optical refractive index, achieving the chip-scale AOM.
@@
The drawback of using piezoelectric materials as the optical waveguide is that it requires etching of the piezoelectric film which is still challenging so far, and only moderate optical can be obtained. Moreover, the waveguide cannot be cladded, which is prone to perturbation from the environment. One should also notice that the piezoelectric material also possesses electro-optic effect, such that the strain can induce electric field across the waveguide via the piezoelectricity which in turn changes its refractive index through the electro-optic effect [15]. This may counteract the optomechanical interaction if not designed correctly. To mitigate these problems, a hybrid optical waveguide can be implemented. As mentioned above, the Si and Si3N4 photonic integrated circuits have evolved into a mature platform with low optical losses and compatible with foundry processes [296]. As shown in Fig. 34(d), the waveguide is entirely embedded in the optical cladding (which is usually SiO2), and the piezoelectric film is deposited on top. Depending on the thickness of the top cladding, the piezoelectric layer can be either fully separated from or partially overlapped with the optical mode. However, since the SAW is primarily confined at the surface, it decays exponentially into the substrate, as can be seen in Fig. 34(h). In that sense, the waveguide needs to be brought as close as possible to the surface to have a good acousto-optic overlap. The SAW AOM of the Si MZI has been demonstrated on both AlN [135] and PZT [294] platforms, as shown in Fig. 34(e-f). The modulation of a hybrid LiNbO3 on Si3N4 waveguide has also been achieved recently [297].
@@
The performance of the state-of-the-art SAW AOM is summarized and compared in Table. 5. One can see that released AOMs generally has higher efficiency (smaller ) than unreleased structures. However, the unreleased AOM has the advantages of easier fabrication, better thermal conductivity, and high-power handling capability. Although the photonic crystal structure demonstrates the smallest due to its tight optical mode confinement, the optical loss is still higher than in other platforms. On the material side, the SAWs made from LiNbO3 show a higher SAW excitation efficiency due to its high electromechanical coupling coefficient. In the future, can be further reduced with the help of novel structures and new materials, targeting at higher microwave to acoustic wave conversion efficiency, smaller acoustic losses, and a larger optomechanical coupling rate (which benefits from a smaller mechanical and optical mode volume, better mode overlap, and larger photoelastic coefficient).
@@ 5.2 Bulk Acoustic Wave Based AOM
Apart from the surface acoustic wave, a bulk acoustic wave (BAW) that transmits vertically in the substrate has also been utilized in modulating the Si3N4 microring resonator [12]. As shown in Fig. 36(a, b), the device shares the same structure as the one shown in Fig. 7, where a disk-shaped actuator is designed to tune the optical resonance of a Si3N4 resonator. If we drive the actuator at the microwave frequency in stead, an acoustic wave will be excited vertically and confined in an acoustic Fabry-Pérot cavity, forming the so-called High-overtone Bulk Acoustic wave Resonator (HBAR) mode. The mode distribution of one typical HBAR mode is illustrated in Fig. 36(c). We can see that the acoustic wave is confined laterally beneath the actuator and is distributed uniformly throughout the substrate. Since the HBAR shown here is a longitudinal wave, the stress is dominated by the vertical component. On the other hand, transverse HBAR with large shear stress has also been demonstrated on LiNbO3 [310, 311]. An efficient AOM can be achieved by placing the optical waveguide on the antinode of the stress distribution.
The HBAR modes can be found from its electromechanical S11 response as presented in Fig. 37(a) and (d). Similarly to an optical Fabry-Pérot cavity, there exists a series of equally spaced mechanical resonances, covering multiple octaves in the microwave frequencies. The mode spacing of 17.5 MHz is primarily determined by the thickness of the substrate. The resonances have a linewidth of 3 MHz, corresponding to a mechanical of 1000. In addition, the envelope of these resonances is periodically modulated. This is due to the presence of acoustic resonance in the SiO2 layer. The coupling between the SiO2 and Si cavities further modifies the mechanical dispersion and perturb the mode spacing. This behavior can be captured by the Mason model, as detailed in Ref. [12].
Each of the acoustic resonances can modulate the optical resonator as manifested in the optomechanical S21 measurement in Fig. 37(b, c) for transverse electric (TE) and transverse magnetic (TM) optical modes. Different from S11 where the resonance dips are less than 1 dB, the S21 shows clear peaks with high contrast ( dB) between resonance and anti-resonance due to the high sensitivity and large signal to noise ratio (SNR) of the optical readout. Maximum S21 of -65 dB can be obtained for both TE and TM modes which leads to a of 200 V according to Eq.28. This is much higher than the SAW AOM.
@@
Although the BAW AOM demonstrated so far has lower efficiency than the SAW AOM, it possesses advantages that are preferred in specific applications. First, since the HBAR mode is distributed uniformly in the vertical direction of the substrate, the optical waveguide can be buried deep inside the cladding to maintain a high optical and immunity to the environment perturbation. Also, the out-of-plane transmission of the acoustic wave makes it possible for independent optimization of the in-plane optical circuits. Secondly, the acoustic wave is confined laterally by the actuator, which guarantees a compact size and a low cross-talk between adjacent actuators [314]. Lastly, since the acoustic resonance is mainly dictated by the thickness of each layer, the lateral degree of freedom is released such that larger critical dimensions of the actuator can be designed, in contrast to the sub-micron electrodes of SAW IDT. This not only eases the fabrication, but also increases the power handling capability. On the other hand, since the thickness of each layer can be well controlled over the whole wafer, HBAR actuators with the same frequency can be massively produced, which would lower the cost. In the future, the AOM efficiency can be improved by engineering the mechanical and optical mode, such as coupling HBAR with a photonic crystal. In addition, materials with larger piezoelectric coefficient, such as LiNbO3 and BaTiO3, will be of great interest for excitation of HBAR modes [310].

### v2

@@ 4.1.2 Wavelength Switchable RSOA-Based External Cavity Laser
There are two frequency tuning schemes supported by the Vernier filter. The first one is wavelength switching by tuning the resonance alignment of the two resonators from one pair to another that is one FSR away, as illustrated in Fig. 25(e). This frequency tuning is discrete but can cover a large range (determined by the FSR) by only tuning a small amount of the resonant frequency of one microring (determined by the FSR difference), which is preferable for piezoelectric actuators. In the experiment, a heater is necessary to initially align the two resonators which are generally far off from each other due to the fabrication variability. Following that, the resonance of one resonator can be tuned by its piezoelectric actuator relative to the other. The wavelength switching direction is determined by which microring is actuated. Figure 25(f) shows the emission wavelength under two switching states that are separated by an FSR. The switching speed is tested by applying 8 V 100 kHz square wave and a 7 ns rise time is achieved. The second scheme is the continuous frequency tuning by actuating the two microrings concurrently. As illustrated in Fig. 25(g), a frequency excursion of 1.5 GHz is generated under 100 kHz triangular wave with 11 V. To suppress the nonlinearity, two algorithms are combined to correct the input voltage, including the frequency response inversion [270] and interative signal correction. After linearization, the RMS nonlinearity changes from 15 MHz to 3.8 MHz, as shown in Fig. 25(h).
@@ 4.1.3 High Power and Tunable Extended-DBR Laser
Table 4 summarizes the performances of the piezoelectric tuned hybrid lasers and compares them with the state-of-the-art hybrid or heterogeneous integrated lasers with fixed frequency, as well as lasers that are tuned by the heater or electro-optic effect. We can see most of the lasers with sub-kHz linewidth are achieved by self-injection locking a DFB laser or an ECL with E-DBR to an ultra-high Si3N4 resonator. The frequency noise is primarily limited by the thermal-TRN of the Si3N4 [266]. A recent work further suppresses noise beyond TRN by locking an E-DBR SIL laser to a Fabry-Pérot cavity via the PDH technique [264]. While the PDH signal was generated by an off-chip electro-optic modulator, the required phase modulation may also be generated via a piezoelectric actuator on chip signal has been generated on-chip by the piezoelectric actuator that is compatible with the Si3N4 PICs [54]. The thermo-optic tuning has achieved the largest tuning range at the expense of high power consumption and slow speed, and the tuning range decreases as the speed increases. On the other hand, electro-optic tuning based on LiNbO3 PICs presents the highest speed up to 600 MHz. However, the laser linewidth is above kHz even with SIL. A recent study reveals that the charge noise of the LiNbO3 ultimately sets the noise floor of the hybrid laser [281]. As for the piezoelectric tuning, lasers with sub-kHz linewidth, over 1 GHz tuning range, and hundreds of kHz modulation speed have been demonstrated. So far, these experiments are implemented by edge coupling the DFB laser or gain chip to the Si3N4 photonic chip. In the furture, the development heterogeneous integration of the laser with piezoelectric actuators and the photonic wirebonding technique will be of great interest for building a more compact and stable chip-scale laser system [259, 263].
@@ 4.2.1 Voltage Initiation and Tuning of Soliton Microcombs
4.2.2 Fast Soliton Microcomb Actuation and Locking
@@
The above experiment only demonstrates the stabilization of one comb line that is involved in the feedback loop. To show the stability of the entire comb, the out-of-loop comb needs to be characterized. To measure the out-of-loop beat signal and its phase noise, a modified setup is used as shown in Fig. 29(a). The pump laser’s frequency to generate the soliton microcomb is shifted by 77.0 MHz via a fibre-coupled AOM. The reason to shift the microcomb’s pump frequency is to cancel out the drift and noise caused by the imbalanced paths in delayed self-homodyne measurement, by detecting the beatnote (77.0 MHz shift) between the pump lines of the microcomb and the EO comb. By down-mixing the 77.0 MHz heterodyne beatnote signal using the same microwave source that drives the AOM, the feedback signal is applied to the laser current such that the pump laser’s frequency is stabilized and the noise in the delayed self-homodyne measurement is removed.
@@
In addition to the quasi-DC tuning, the HBAR modes of the AlN actuator provide a novel functionality for error signal generation using the PDH technique. Figure 31(a) shows the experimental setup to generate PDH error signals using HBAR modes. The measured response of the AlN actuation, up to 400 MHz, is plotted in the linear frequency scale in Fig. 31(b), showing both cases when the laser is on- and off-resonance. Different modulation frequencies corresponding to different HBAR modes are investigated, which are marked with stars in Fig. 31(b). The PDH error signals modulated at these HBAR frequencies are shown in Fig. 31(c), as well as the corresponding microresonator’s resonance. A microwave source providing 8 dBm RF power is used to modulate the Si3N4 microresonator via AlN actuation. The decrease in error signal contrast at higher HBAR frequency is caused by the finite bandwidth of the optical resonance. This PDH signal can be used to stabilize the laser frequency referenced to an on-chip high optical Q Si3N4 microresonator.
4.2.3 Soliton Microcomb Based Parallel LiDAR Engine
The AlN actuator can be a key component for a soliton-based, parallel, FMCW LiDAR [38]. This scheme conventionally relies on scanning the pump laser’s frequency over the soliton existence range , i.e. , with being the resonance frequency, and the boundary of soliton existence detuning range. As a result, the pump frequency chirp is transferred to all soliton comb lines, as shown in Fig. 32(a). The diffractive optics subsequently separates the comb lines spatially, which allows high-speed and parallel acquisition of both velocity and position in each pixel. However, the varying soliton detuning has several limitations. First, the soliton detuning variation leads to changes in soliton spectrum bandwidth and power, which limits the number of usable optical channels. Second, the Raman self-frequency shift [290, 291] causes a varied soliton repetition rate. Third, GHz frequency excursion of the pump laser () requires GHz-wide soliton existence range (), which leads to elevated pump power to several Watts.
@@ 4.2.3 Soliton Microcomb Based Parallel LiDAR Engine
In conclusion, integrated piezoelectric actuators that are linear, fast, non-absorptive, bidirectional, and consume ultralow electric power endow integrated soliton microcombs with capability in power-critical applications, e.g., in space, data centers and portable atomic clocks, in extreme environments such as cryogenic temperatures, or in emerging applications for coherent LiDAR [38], frequency synthesizers [34] and RF photonics [293, 294]. Therefore, this capability is important for tuning and stabilization of soliton microcombs, and may allow soliton injection locking for stabilization and microwave synthesis [287, 295]. While polycrystalline AlN is used in current works, the operation voltage can be reduced by more than 2 times using scandium-doped AlN [296]. In the future, the co-integration of CMOS electronic circuitry on a close-by die paves the way for packaged soliton microcombs with rapid electronic actuation [273].
5 Radio-Frequency Acousto-Optic Modulation
While piezoelectric actuators have found wide applications in tuning of integrated photonic devices, fast actuation at microwave frequencies will excite acoustic waves that enable efficient acousto-optic modulation. At the acoustic resonance, integrated AOMs have been demonstrated with efficiency comparable to or even higher than state-of-the-art electro-optic modulators, albeit at the expense of narrow bandwidth [15, 14, 12, 297, 298]. Nevertheless, it has been applied not only in classical applications, but also in quantum regime working as quantum transducers between microwave and light [201, 299]. In the following, we will review recent progress on chip-scale AOMs that can be classified into three categories in terms of the acoustic mode being utilized: surface acoustic wave, bulk acoustic wave, and optomechanical crystal.
@@ 5.1 Surface Acoustic Wave Based AOM
The surface acoustic wave excited by IDT has been one of the traditional ways of launching acoustic waves on a chip [300], and has become the key technology across a wide range of research areas, such as microfluidics [301], biosensors [302], radio-frequency filters and delay lines [303, 304], and quantum acoustics [305]. As nano-fabrication advances, the ability to create sub-micron electrode fingers has extended the frequency of the acoustic wave to beyond 10 GHz [306, 142]. By interacting with an optical cavity, the stress and strain field of the acoustic wave can be used to modulate the optical refractive index, achieving the chip-scale AOM.
@@
The drawback of using piezoelectric materials as the optical waveguide is that it requires etching of the piezoelectric film which is still challenging for most piezoelectric materials. In recent decade, promising progresses have been made in achieving high-Q optical integrated resonators, where intrinsic Q as high as 3.7 million and 11 million have been demonstrated in AlN [309] and LiNbO3 [310] microring resonators, respectively. Additionally, the waveguide cannot be cladded, which is prone to perturbation from the environment. One should also notice that the piezoelectric material also possesses electro-optic effect, such that the strain can induce electric field across the waveguide via the piezoelectricity which in turn changes its refractive index through the electro-optic effect [15]. This may counteract the optomechanical interaction if not designed correctly. To mitigate these problems, a hybrid optical waveguide can be implemented. As mentioned above, the Si and Si3N4 photonic integrated circuits have evolved into a mature platform which is compatible with foundry processes [311]. Moreover, ultra-high-Q Si3N4 integrated resonators have been demonstrated by multiple groups [312], and intrinsic Q as high as 442 million has been achieved [26]. As shown in Fig. 34(d), the waveguide is entirely embedded in the optical cladding (which is usually SiO2), and the piezoelectric film is deposited on top. Depending on the thickness of the top cladding, the piezoelectric layer can be either fully separated from or partially overlapped with the optical mode. However, since the SAW is primarily confined at the surface, it decays exponentially into the substrate, as can be seen in Fig. 34(h). In that sense, the waveguide needs to be brought as close as possible to the surface to have a good acousto-optic overlap. The SAW AOM of the Si MZI has been demonstrated on both AlN [144] and PZT [307] platforms, as shown in Fig. 34(e-f). The modulation of a hybrid LiNbO3 on Si3N4 waveguide has also been achieved recently [313].
@@
The performance of the state-of-the-art SAW AOM is summarized and compared in Table. 5. One can see that released AOMs generally have higher efficiency (smaller ) than unreleased structures. However, the unreleased AOM has the advantages of easier fabrication, better thermal conductivity, and high-power handling capability. Although the photonic crystal structure demonstrates the smallest due to its tight optical mode confinement, the optical loss is still higher than in other platforms. On the material side, the SAWs made from LiNbO3 show a higher SAW excitation efficiency due to its high electromechanical coupling coefficient. In the future, can be further reduced with the help of novel structures and new materials, targeting at higher microwave to acoustic wave conversion efficiency, smaller acoustic losses, and a larger optomechanical coupling rate (which benefits from a smaller mechanical and optical mode volume, better mode overlap, and larger photoelastic coefficient).
@@ 5.2 Bulk Acoustic Wave Based AOM
Apart from the surface acoustic wave, a bulk acoustic wave that transmits vertically in the substrate has also been utilized in modulating the Si3N4 microring resonator [12]. As shown in Fig. 36(a, b), the device shares the same structure as the one shown in Fig. 7, where a disk-shaped actuator is designed to tune the optical resonance of a Si3N4 resonator. If we drive the actuator at the microwave frequency in stead, an acoustic wave will be excited vertically and confined in an acoustic Fabry-Pérot cavity, forming the so-called HBAR mode. The mode distribution of one typical HBAR mode is illustrated in Fig. 36(c). We can see that the acoustic wave is confined laterally beneath the actuator and is distributed uniformly throughout the substrate. Since the HBAR shown here is a longitudinal wave, the stress is dominated by the vertical component. On the other hand, transverse HBAR with large shear stress has also been demonstrated on LiNbO3 [326, 327]. An efficient AOM can be achieved by placing the optical waveguide on the antinode of the stress distribution.
The HBAR modes can be found from its electromechanical S11 response as presented in Fig. 37(a) and (d). Similarly to an optical Fabry-Pérot cavity, there exists a series of equally spaced mechanical resonances, covering multiple octaves in the microwave frequencies. The mode spacing of 17.5 MHz is primarily determined by the thickness of the substrate. The resonances have a linewidth of 3 MHz, corresponding to a mechanical of 1000. In addition, the envelope of these resonances is periodically modulated. This is due to the presence of an acoustic resonance in the SiO2 layer. The coupling between the SiO2 and Si cavities further modifies the mechanical dispersion and perturbs the mode spacing. This behavior can be captured by the Mason model, as detailed in Ref. [12].
Each of the acoustic resonances can modulate the optical resonator as manifested in the optomechanical S21 measurement in Fig. 37(b, c) for TE00 and TM00 optical modes. Different from S11, where the resonance dips are less than 1 dB, the S21 shows clear peaks with high contrast ( dB) between resonance and anti-resonance due to the high sensitivity and large signal to noise ratio (SNR) of the optical readout. Maximum S21 of -65 dB can be obtained for both TE00 and TM00 modes which leads to a of 200 V according to Eq.29. This is much higher than the SAW AOM.
@@
Although the BAW AOM demonstrated so far has lower efficiency than the SAW AOM, it possesses advantages that are preferred in specific applications. First, since the HBAR mode is distributed uniformly in the vertical direction of the substrate, the optical waveguide can be buried deep inside the cladding to maintain a high optical and immunity to the environment perturbation. Also, the out-of-plane transmission of the acoustic wave makes it possible for independent optimization of the in-plane optical circuits. Secondly, the acoustic wave is confined laterally by the actuator, which guarantees a compact size and a low cross-talk between adjacent actuators [330]. Lastly, since the acoustic resonance is mainly dictated by the thickness of each layer, the lateral degree of freedom is released such that larger critical dimensions of the actuator can be designed, in contrast to the sub-micron electrodes of SAW IDT. This not only eases the fabrication, but also increases the power handling capability. On the other hand, since the thickness of each layer can be well controlled over the whole wafer, HBAR actuators with the same frequency can be massively produced at the wafer scale, which would improve the yield and lower the cost. In the future, the AOM efficiency can be improved by engineering the mechanical and optical mode, such as coupling HBAR with a photonic crystal. In addition, materials with larger piezoelectric coefficient, such as LiNbO3 and BaTiO3, will be of great interest for excitation of HBAR modes [326].